OpenStudy (anonymous):

find all numbers that must be excluded from the domain of the rational expression 9/x−6

jimthompson5910 (jim_thompson5910):

You cannot divide by zero. So x-6 cannot be equal to zero. So this means that x can't be equal to 6 (since x-6=6-6=0). So the domain is the set of all real numbers but x can't equal 6.

OpenStudy (anonymous):

So how do i show work for that?

jimthompson5910 (jim_thompson5910):

show that solving for x-6=0 gives you x=6. So basically if x=6, then x-6=0. This then would mean that you have to exclude x=6 from the domain.